2.5kg of water is heated in an open container for 5 minutes using a heater of power 2KW. Which of the following statements about the temperature of the water is true?(the SHC of water= 4200Jkg-1k-1)

A. The change in temperature of the water is 58.30C

B. The change in temperature of the water is 57.30C

C. The change in temperature of the water is greater 57.30C

D.If the initial temperature of the water was 23.70C then its final temperature would be 8100C


1
Expert's answer
2021-11-03T10:23:26-0400

The amount of heat required to change the temperature of the water can be found as follows:


"Q=mc\Delta T."

From the other hand, we can write:


"Q=Pt."

Equating these two expressions, we get:


"mc\Delta T=Pt,""\Delta T=\dfrac{Pt}{mc},""\Delta T=\dfrac{2.0\times10^3\ W\times5\ min\times\dfrac{60\ s}{1\ min}}{2.5\ kg\times4200\ \dfrac{J}{kg\!\times\!^{\circ}C}}=57.3^{\circ}C."

As we can see from the calculations, the change in temperature of the water is "57.3^{\circ}C". Therefore, the correct answer is (B).

Answer:

B. The change in temperature of the water is 57.30C